# Stable cohomology of universal character varieties

### v1

@@ 1. Introduction
The degree $k$ cohomology group of the moduli space of curves of genus $g$ is independent of $g$ for $k\leq\frac{2g-2}{3}$. These stable cohomology groups were computed by Madsen and Weiss [MW07], confirming a conjecture of Mumford [Mum83]. Subsequently, Ellenberg, Venkatesh, and Westerland initiated an analogous study of cohomological stability for Hurwitz spaces, which parametrize branched $G$-covers of $\mathbb{A}^{1}$ for a finite group $G$ or, equivalently, $G$-local systems on $\mathbb{A}^{1}$ minus finitely many branch points [EVW16]. More recently, Landesman and Levy computed the stable cohomology groups of these $G$-Hurwitz spaces [LL25] as the number of branch points goes to $\infty$, and Putman [Put23, Theorem C] proved a homological stability result for $G$-Hurwitz spaces of a genus $g$ curve with one boundary component in the limit as $g\to\infty$.
@@
parametrizing $\operatorname{PGL}_{n}(\mathbb{C})$-local systems on $C$ via the monodromy representation. This character variety breaks into pieces indexed by $\mathbb{Z}/n\mathbb{Z}$, based on the obstruction to lifting a homomorphism to $\operatorname{PGL}_{n}(\mathbb{C})$ to $\operatorname{GL}_{n}(\mathbb{C})$, which is locally constant. We study the stable cohomology of the components corresponding to $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$, which are smooth. See §2 for precise definitions and further details.
@@
be the first rational cohomology group of a closed orientable surface of genus $g$, viewed as a $\operatorname{Mod}_{g}$-module or, equivalently, a local system on $\mathcal{M}_{g}$. We consider $V_{g}$ as a graded object, supported in degree $1$, so $V_{g}[-i]$ is supported in degree $i+1$. We also define
@@
Corollary 1.2 follows from Theorem 1.1 via Looijenga’s results on the cohomology of $\mathcal{M}_{g}$ with symplectic coefficients [Loo96]. See Section 4.2 for further details on the stable cohomology groups. The dimensions of the stable cohomology groups of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{M}_{g},n,1)$ for $n\leq 7$ in even degrees up to 20 are displayed in Figure 1.
@@
We prove this degeneration using the relative nonabelian Hodge correspondence [Sim94], purity of the cohomology of the moduli space of Higgs bundles [HRV15, Corollary 1.3.2], and a splitting result for pure Hodge modules [Sai90, (4.5.4)]. We also present a second proof using only classical Hodge theory and manipulations in the derived category of constructible sheaves. Theorem 1.1 is deduced from the degeneration of the Leray-Serre spectral sequence for $\pi$ using previously known results on generators and relations of the cohomology of the fibers ([Mar02, Theorem 7] and [HRV08, Lemma 4.1.12]).
We also consider universal $G$-character varieties over $\mathcal{M}_{g,1}$, for $G$ = $\operatorname{GL}_{n}$ or $\operatorname{SL}_{n}$. The fiber of the universal $G$-character variety over a punctured curve $C$ is given by
@@
For $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ and $k\leq 2(g-1)(n-1)+2$, the cohomology group $H^{k}(M^{\operatorname{SL}}_{B}(\mathcal{C}_{g,1},n,d))$ is isomorphic to the degree $k$ part of the graded vector space
@@
Similarly, for $k\leq 2(g-1)(n-1)+2$, the cohomology group $H^{k}(M^{\operatorname{GL}}_{B}(\mathcal{C}_{g,1},n,d))$ is isomorphic to the degree $k$ part of the graded vector space
@@
We are not aware of a relative non-abelian Hodge correspondence for $\operatorname{SL}$- and $\operatorname{GL}$-character varieties in the literature that would allow a direct proof of Theorem 1.4. Instead, we deduce this statement from the above results for $\operatorname{PGL}$-character varieties. The final statements about stable cohomology in low degrees follows, just as Corollary 1.2 follows from Theorem 3.1, using results of Kawazumi and Randal-Williams on the cohomology of $\mathcal{M}_{g,1}$ with twisted coefficients [Kaw08, RW18]. See Section 4.3 for further discussion of the stable cohomology. The dimensions of the stable cohomology groups of $M_{B}^{\operatorname{SL}}(\mathcal{C}_{g,1},2,1)$ in even degrees up to 20 are displayed in Figure 1.
@@
By Theorem 1.1, the rational cohomology of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ is independent of the choice of $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$. We do not know any way of deducing this from first principles. The fibers of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ are naturally viewed as algebraic varieties defined over $\mathbb{Q}[e^{2\pi i/n}]$. The fibers for varying $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ are Galois conjugate, giving an isomorphism of cohomology (with $\mathbb{Q}_{\ell}$-coefficients). Such arguments do not apply to the total space of the universal character varieties, which are constructed analytically (Definition 2.1). Our proof of Theorem 1.1 depends on the relative non-abelian Hodge correspondence, which produces a homeomorphism over $\mathcal{M}_{g}$ between the universal character variety $M_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ on the Betti side and a universal moduli space of semistable Higgs bundles of degree $d$ $M_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ on the Dolbeault side. The fact that the cohomology is independent of the choice of $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ does not have an obvious geometric explanation in the Higgs bundle setting, though the result fits into a more general conjecture about the behavior of Gopakumar-Vafa invariants on Calabi-Yau 3-folds. See [MS23, §0.3].
@@ 2. Universal character varieties
Let $\mu_{n}\subset\mathbb{C}$ be the $n$th roots of unity, which we identify with $\mathbb{Z}/n\mathbb{Z}$ via $d\mapsto e^{2\pi id/n}$. By the short exact sequence
@@
Assume $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$. Then $X_{g,n,d}$ is smooth and irreducible, and $\operatorname{PGL}_{n}(\mathbb{C})$ acts with finite stabilizers [HRV08, Theorem 2.2.12 and Corollary 3.5.5]. In particular, $\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)$ is a smooth orbifold, and projection to the coarse space $\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)\to M_{B}^{\operatorname{PGL}}(S_{g},n,d)$ induces an isomorphism
@@
Although the base and fiber of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ each carry a natural algebraic structure, our construction does not immediately endow the total space with such a structure. In order to apply techniques from algebraic geometry to study the rational cohomology of $M_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$, we pass to a homeomorphic algebraic moduli space of semistable Higgs bundles on the Dolbeault side of the nonabelian Hodge correspondence.
@@
on which $G(\mathbb{C})$ acts by conjugation. If $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ then $Y^{G}_{g,n,d}$ is smooth and irreducible, and the $G(\mathbb{C})$-action factors through a free action of $\operatorname{PGL}_{n}(\mathbb{C})$ [HRV08, Corollary 2.2.7]. In these cases, the degree $d$ $G$-character variety is the quotient
@@
Because the outer action of the mapping class group $\operatorname{Mod}_{g,1}\coloneq\pi_{0}(\operatorname{Diff}^{+}(S_{g,1}))$ on $\pi_{1}(S_{g,1})$ preserves the conjugacy class of $\gamma$ [FM12, Theorem 8.8], $\mathcal{M}_{B}^{G}(S_{g,1},n,d)$ inherits an action of the mapping class group $\operatorname{Mod}_{g,1}$ of $S_{g,1}$, by precomposition. Let $\mathcal{T}_{g,1}$ denote the Teichmüller space of $S_{g,1}$, which likewise carries a natural action of $\operatorname{Mod}_{g,1}$.
@@
When $d=0$, the condition $\phi(\gamma)=\operatorname{Id}_{n}$ from (6) is equivalent to the condition that $\phi$ factors through the fundamental group of the unpunctured surface $\pi_{1}(S_{g,1})\to\pi_{1}(S_{g}).$ The outer action of $\operatorname{Mod}_{g,1}$ on $\pi_{1}(S_{g})$ then factors through the outer action of $\operatorname{Mod}_{g}$, and $M_{B}^{G}(\mathcal{C}_{g,1},n,0)$ is the pullback of an analogously defined universal degree 0 $G$-character variety over $\mathcal{M}_{g}$. Hence, one can study degree 0 relative $G$-character varieties for families of unpunctured curves, as in [Sim94, §6 and §9].
However, if $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ then $M_{B}^{G}(\mathcal{C}_{g,1},n,d)$ is not the pullback of a space over $\mathcal{M}_{g}$, because the kernel of the map $\pi_{1}(\mathcal{M}_{g,1})\to\pi_{1}(\mathcal{M}_{g})$, acts nontrivially on $M_{B}^{G}(\mathcal{C}_{g,1},n,d)$. This can be seen from the description of its outer action on $\pi_{1}(S_{g})$ in terms of Dehn twists [FM12, Fact 4.7].
@@
We now discuss relations among the universal $G$-character varieties for varying groups $G$. Note that $\Gamma:=\mathrm{Hom}(\pi_{1}(S_{g,1}),\mu_{n})$ acts on $Y_{g,n,d}^{G}$ via the inclusion of $\mu_{n}$ in the center of $G$, and this action descends to an action on the character variety $M_{B}^{G}(S_{g,1},n,d)$. Passing to the universal setting, the relative group object
@@ 3. Degeneration of the Leray spectral sequence
The first statement in Theorem 3.1 is a universal form of the degeneration of the Leray spectral sequence in the following sense: an object $A$ in a bounded derived category $D^{b}(\mathcal{A})$ splits as the direct sum of its cohomology sheaves if and only if for any cohomological functor $F$ from $D^{b}(\mathcal{A})$ to an abelian category, the corresponding spectral sequence $E_{2}^{pq}=F(\mathcal{H}^{q}(A)[p])\implies F(A[p+q])$ degenerates at $E_{2}$ [Del68, Proposition 1.2].
@@
One may also give a third proof of Theorem 3.1, similar in spirit to the two presented here, but different in technical detail, via spreading out over a finitely generated $\mathbb{Z}$-algebra, reducing modulo large finite primes, and applying the theory of weights in the $\ell$-adic cohomology of local systems on varieties over finite fields, along the lines of [BBD82, Theorem 5.4.5]. Like the two proofs presented here, this argument relies on the theory of weights in the cohomology of local systems on algebraic varieties [Del75]. However, to present this third proof in detail would take us too far from the setting of this paper, which is rooted in geometry over the complex numbers.
@@ 3.1. The moduli spaces
A Higgs bundle $(E,\theta)$ is stable (resp. semistable) if $\mu(F)<\mu(E)$ (resp. $\mu(F)\leq\mu(E))$ for every $\theta$-invariant subbundle $F\subset E$. When $d$ is coprime to $n$, a Higgs bundle is semistable if and only if it is stable. Let $M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)$ denote the coarse moduli space of semistable Higgs bundles of degree $d$ and rank $n$ on $C$, as constructed by Simpson [Sim94].
For moduli of $\operatorname{SL}$- and $\operatorname{PGL}$-Higgs bundles, we follow the definitions from [HT03, §2]. Fixing an auxiliary degree $d$ line bundle $L$ on $C$, the moduli space of $\operatorname{SL}$-Higgs bundles $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)$ is the subspace of $M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)$ parametrizing Higgs bundles $(E,\theta)$ such that $\det E\cong L$ and $\operatorname{tr}\theta=0$, when $\theta$ is viewed as an $\omega_{C}$-twisted endomorphism of $E$.
@@
Here, the group of $n$-torsion points $\mathrm{Jac}(C)[n]$ acts on $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)$ by $L\cdot(E,\theta):=(L\otimes E,\theta).$ For a different choice of degree $d$ line bundle $L^{\prime}$, there is a non-canonical isomorphism $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)\cong M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L^{\prime})$ given by tensoring with an $n^{\mathrm{th}}$ root of $L^{-1}\otimes L^{\prime}\in\mathrm{Jac}(C)$. It follows that $M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)$ is well-defined by (9) and independent of $L$.
These Higgs moduli spaces are related to the character varieties on the Betti side of the nonabelian Hodge correspondence, as follows [Sim94].
Nonabelian Hodge Correspondence.
@@
$M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)\cong M^{\operatorname{GL}}_{B}(S_{g,1},n,d),\quad\mbox{ and }\quad M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,d)\cong M^{\operatorname{SL}}_{B}(S_{g,1},n,d).$

### v2

@@ 1. Introduction
The degree $k$ cohomology group of the moduli space of curves of genus $g$ is independent of $g$ for $k\leq\frac{2g-2}{3}$. These stable cohomology groups were computed by Madsen and Weiss [31], confirming a conjecture of Mumford [39]. Subsequently, Ellenberg, Venkatesh, and Westerland initiated an analogous study of cohomological stability for Hurwitz spaces, which parametrize branched $G$-covers of $\mathbb{A}^{1}$ for a finite group $G$ or, equivalently, $G$-local systems on $\mathbb{A}^{1}$ minus finitely many branch points [14]. More recently, Landesman and Levy computed the stable cohomology groups of these $G$-Hurwitz spaces [29] as the number of branch points goes to $\infty$, and Putman [43, Theorem C] proved a homological stability result for $G$-Hurwitz spaces of a genus $g$ curve with one boundary component in the limit as $g\to\infty$.
@@
parametrizing $\operatorname{PGL}_{n}(\mathbb{C})$-local systems on $C$ via the monodromy representation. This character variety breaks into pieces indexed by $\mathbb{Z}/n\mathbb{Z}$, based on the obstruction to lifting a homomorphism to $\operatorname{PGL}_{n}(\mathbb{C})$ to $\operatorname{GL}_{n}(\mathbb{C})$, which is locally constant. We study the stable cohomology of the components $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ corresponding to $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$, which are smooth; the appendix explains why this is the same as stable intersection cohomology for any $d\in(\mathbb{Z}/n\mathbb{Z})$. See §2 for precise definitions and further details.
@@
be the first rational cohomology group of a closed orientable surface of genus $g$, viewed as a $\operatorname{Mod}_{g}$-module or, equivalently, a local system on $\mathcal{M}_{g}$. We consider $V_{g}$ as a graded object, supported in degree $1$, so $V_{g}[-i]$ is supported in degree $i+1$. By convention, the grading on the cohomology of a graded local system is given by the sum of the cohomological degree and the degree of the local system, e.g. $H^{k}(\mathcal{M}_{g};V_{g}[-2])$ is of degree $k+3$. We also define
@@
Corollary 1.2 follows from Theorem 1.1 via Looijenga’s results on the cohomology of $\mathcal{M}_{g}$ with symplectic coefficients [30]. See Section 4.2 for further details on the stable cohomology groups. The dimensions of the stable cohomology groups of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,1)$ for $n\leq 7$ in even degrees up to 20 are displayed in Figure 1.
@@
We prove this degeneration using the relative nonabelian Hodge correspondence [49], purity of the cohomology of the moduli space of Higgs bundles [18, Corollary 1.3.2], and a splitting result for pure Hodge modules [45, (4.5.4)]. We also present a second proof using only classical Hodge theory and manipulations in the derived category of constructible sheaves. Theorem 1.1 is deduced from the degeneration of the Leray-Serre spectral sequence for $\pi$ using previously known results on generators and relations of the cohomology of the fibers ([32, Theorem 7] and [19, Lemma 4.1.12]).
We also consider universal $G$-character varieties over $\mathcal{M}_{g,1}$, for $G$ = $\operatorname{GL}_{n}$ or $\operatorname{SL}_{n}$. The fiber of the universal $G$-character variety $M^{G}_{B}(\mathcal{C}_{g,1},n,d)$ over a punctured curve $C$ is given by
@@
For $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ and $k\leq 2(g-1)(n-1)+2$, the cohomology group $H^{k}(M^{\operatorname{SL}}_{B}(\mathcal{C}_{g,1},n,d);\mathbb{Q})$ is isomorphic to the degree $k$ part of the graded vector space
@@
Similarly, for $k\leq 2(g-1)(n-1)+2$, the cohomology group $H^{k}(M^{\operatorname{GL}}_{B}(\mathcal{C}_{g,1},n,d);\mathbb{Q})$ is isomorphic to the degree $k$ part of the graded vector space
@@
The final statements about stable cohomology in low degrees follows, just as Corollary 1.2 follows from Theorem 3.1, using results of Kawazumi and Randal-Williams on the cohomology of $\mathcal{M}_{g,1}$ with twisted coefficients [26, 44]. See Section 4.3 for further discussion of the stable cohomology. The dimensions of the stable cohomology groups of $M_{B}^{\operatorname{SL}}(\mathcal{C}_{g,1},2,1)$ in even degrees up to 20 are displayed in Figure 1.
@@
Although we initially hoped that each universal degree $d$ coprime to $n$ character variety could naturally be defined over $\mathcal{M}_{g}$ instead of $\mathcal{M}_{g,1}$, we no longer believe that this is the case for $\operatorname{GL}$ and $\operatorname{SL}$; see Remark 2.4. This is also consistent with our belief that the most natural path to proving a relative nonabelian Hodge correspondence in coprime degree would be to reinterpret the question in terms of terms of degree 0 filtered objects on a punctured curve, extending Simpson’s proof of a fiberwise bijection [48]. However, to our knowledge such a result is not in the literature, and so we reduce to using Simpson’s degree 0 relative nonabelian Hodge result (for $\operatorname{PGL}$) [49].
Remark 1.6.
By Theorem 1.1, the rational cohomology of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ is independent of the choice of $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$. We do not know any way of deducing this from first principles. The fibers of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ are naturally viewed as algebraic varieties defined over $\mathbb{Q}[e^{2\pi i/n}]$. The fibers for varying $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ are Galois conjugate, giving an isomorphism of cohomology (with $\mathbb{Q}_{\ell}$-coefficients). Such arguments do not apply to the total space of the universal character varieties, which are constructed analytically (Definition 2.1). Our proof of Theorem 1.1 depends on the relative non-abelian Hodge correspondence, which produces a homeomorphism over $\mathcal{M}_{g}$ between the universal character variety $M_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ on the Betti side and a universal moduli space of semistable Higgs bundles of degree $d$ $M_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ on the Dolbeault side. The fact that the cohomology is independent of the choice of $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ does not have an obvious geometric explanation in the Higgs bundle setting, though the result fits into a more general conjecture about the behavior of Gopakumar-Vafa invariants on Calabi-Yau 3-folds. See [34, §0.3].
@@ 2. Universal character varieties
Let $\mu_{n}\subset\mathbb{C}$ be the group of $n$th roots of unity, which we identify with $\mathbb{Z}/n\mathbb{Z}$ via $d\mapsto e^{2\pi id/n}$. By the short exact sequence
@@
Assume $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$. Then $X_{g,n,d}$ is smooth and irreducible, and $\operatorname{PGL}_{n}(\mathbb{C})$ acts with finite stabilizers [19, Theorem 2.2.12 and Corollary 3.5.5]. In particular, $\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)$ is a smooth orbifold, and projection to the coarse space $\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)\to M_{B}^{\operatorname{PGL}}(S_{g},n,d)$ induces an isomorphism
@@
Although the base and fiber of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ each carry a natural algebraic structure, our construction does not immediately endow the total space with such a structure. In order to apply techniques from algebraic geometry to study the rational cohomology of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$, we pass to a homeomorphic algebraic moduli space of semistable Higgs bundles on the Dolbeault side of the nonabelian Hodge correspondence.
@@
on which $G(\mathbb{C})$ acts by conjugation. If $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ then $Y^{G}_{g,n,d}$ is smooth and irreducible, and the $G(\mathbb{C})$-action factors through a free action of $\operatorname{PGL}_{n}(\mathbb{C})$ [19, Corollary 2.2.7]. In these cases, the degree $d$ $G$-character variety is the quotient
@@
Because the outer action of the mapping class group $\operatorname{Mod}_{g,1}\coloneq\pi_{0}(\operatorname{Diff}^{+}(S_{g,1}))$ on $\pi_{1}(S_{g,1})$ preserves the conjugacy class of $\gamma$ [15, Theorem 8.8], $\mathcal{M}_{B}^{G}(S_{g,1},n,d)$ inherits an action of the mapping class group $\operatorname{Mod}_{g,1}$ of $S_{g,1}$, by precomposition. Let $\mathcal{T}_{g,1}$ denote the Teichmüller space of $S_{g,1}$, which likewise carries a natural action of $\operatorname{Mod}_{g,1}$.
@@
When $d=0$, the condition $\phi(\gamma)=\operatorname{Id}_{n}$ from (6) is equivalent to the condition that $\phi$ factors through the fundamental group of the unpunctured surface $\pi_{1}(S_{g,1})\to\pi_{1}(S_{g}).$ The outer action of $\operatorname{Mod}_{g,1}$ on $\pi_{1}(S_{g})$ then factors through the outer action of $\operatorname{Mod}_{g}$, and $M_{B}^{G}(\mathcal{C}_{g,1},n,0)$ is the pullback of an analogously defined universal degree 0 $G$-character variety over $\mathcal{M}_{g}$. Hence, one can study degree 0 relative $G$-character varieties for families of unpunctured curves, as in [49, §6 and §9].
However, if $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ then $M_{B}^{G}(\mathcal{C}_{g,1},n,d)$ is not the pullback of a space over $\mathcal{M}_{g}$, because the kernel of the map $\pi_{1}(\mathcal{M}_{g,1})\to\pi_{1}(\mathcal{M}_{g})$, acts nontrivially on $M_{B}^{G}(\mathcal{C}_{g,1},n,d)$. This can be seen from the description of its outer action on $\pi_{1}(S_{g})$ in terms of Dehn twists [15, Fact 4.7].
@@
We now discuss relations among the universal $G$-character varieties for varying groups $G$. Note that $\Gamma:=\mathrm{Hom}(\pi_{1}(S_{g,1}),\mu_{n})$ acts on $Y_{g,n,d}^{G}$ by multiplication, via the inclusion of $\mu_{n}$ in the center of $G$, and this action descends to an action on the character variety $M_{B}^{G}(S_{g,1},n,d)$. Passing to the universal setting, the relative group object
@@ 3. Degeneration of the Leray spectral sequence
The first statement in Theorem 3.1 is a universal form of the degeneration of the Leray spectral sequence in the following sense: an object $A$ in a bounded derived category $D^{b}(\mathcal{A})$ splits as the direct sum of its cohomology sheaves if and only if for any cohomological functor $F$ from $D^{b}(\mathcal{A})$ to an abelian category, the corresponding spectral sequence $E_{2}^{pq}=F(\mathcal{H}^{q}(A)[p])\implies F(A[p+q])$ degenerates at $E_{2}$ [9, Proposition 1.2].
@@
One may also give a third proof of Theorem 3.1, similar in spirit to the two presented here, but different in technical detail, via spreading out over a finitely generated $\mathbb{Z}$-algebra, reducing modulo large finite primes, and applying the theory of weights in the $\ell$-adic cohomology of local systems on varieties over finite fields, along the lines of [2, Theorem 5.4.5]. Like the two proofs presented here, this argument relies on the theory of weights in the cohomology of local systems on algebraic varieties [12]. However, to present this third proof in detail would take us too far from the setting of this paper, which is rooted in geometry over the complex numbers.
@@ 3.1. The moduli spaces
A Higgs bundle $(E,\theta)$ is stable (resp. semistable) if $\mu(F)<\mu(E)$ (resp. $\mu(F)\leq\mu(E))$ for every $\theta$-invariant subbundle $F\subset E$. When the degree and rank are coprime, a Higgs bundle is semistable if and only if it is stable. Let $M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)$ denote the coarse moduli space of semistable Higgs bundles of degree $d$ and rank $n$ on $C$, as constructed by Simpson [49].
For moduli of $\operatorname{SL}$- and $\operatorname{PGL}$-Higgs bundles, we follow the definitions from [20, §2]. Fixing an auxiliary degree $d$ line bundle $L$ on $C$, the moduli space of $\operatorname{SL}$-Higgs bundles $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)$ is the subspace of $M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)$ parametrizing Higgs bundles $(E,\theta)$ such that $\det E\cong L$ and $\operatorname{tr}\theta=0$, when $\theta$ is viewed as an $\omega_{C}$-twisted endomorphism of $E$.
@@
Here, the group of $n$-torsion points $\mathrm{Jac}(C)[n]$ acts on $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)$ by $M\cdot(E,\theta):=(M\otimes E,\theta).$ For a different choice of degree $d$ line bundle $L^{\prime}$, there is a non-canonical isomorphism $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)\cong M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L^{\prime})$ given by tensoring with an $n^{\mathrm{th}}$ root of $L^{-1}\otimes L^{\prime}\in\mathrm{Jac}(C)$. It follows that $M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)$ is well-defined by (9) and independent of $L$.
These Higgs moduli spaces are related to the character varieties on the Betti side of the nonabelian Hodge correspondence, as follows [22, Theorem 4.3].
Fiberwise Nonabelian Hodge Correspondence.
@@
$M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)\cong M^{\operatorname{GL}}_{B}(S_{g,1},n,d),\quad\mbox{ and }\quad M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)\cong M^{\operatorname{SL}}_{B}(S_{g,1},n,\deg L).$